### v1

@@ 8.1. Case $d\geq 2$, the lower bound
$\tau^{-cj}\ll(j\log\tau)^{-2/d}\leq r_{j}^{d+1}\ll y_{j},$
where the last inequality relies on the assumption $\inf J\ggg_{\gamma_{2}}1$ and (8.12). Hence $\tau^{-ck}\ll y_{j}\tau^{-c(k-j)}$, so $\sum\nolimits_{j,k\in J_{1},j\leq k}\tau^{-ck}\ll y_{J_{1}}$ as for the first sum. The fourth sum can be handled similarly to the third one. As $\sum_{J_{1}}y_{j}\leq(1+y_{J_{1}})^{3/2}$, we have justified the upper bound for $J_{1}$.
@@
Set $m:=|J_{2}|$. We start with the case where $m$ is very small compared to $n=|J_{2}\sqcup J_{3}|$, more precisely we assume $m^{2}\leq n$. In this scenario, we have by the Cauchy-Schwarz inequality and Lemma˜8.11,
@@
(8.20) $\mathbb{E}\bigl{[}Z_{J^{\prime}_{2}}^{2}\bigr{]}$ $=\sum_{|j-k|<\sqrt{m}}\mathbb{E}[Z_{j}Z_{k}]+\sum_{|j-k|\geq\sqrt{m}}\mathbb{E}[Z_{j}Z_{k}]$
(8.21) $\ll\sqrt{m}\sum_{j\in J^{\prime}_{2}}\mathbb{E}[Z_{j}^{2}]+\Bigl{\lvert}\sum_{|j-k|\geq\sqrt{m}}\mathbb{E}[Z_{j}Z_{k}]\Bigr{\rvert}$
@@
For subscripts $j\leq k\in J_{2}^{\prime}$ such that $|j-k|\geq\sqrt{m}$, we have by (8.17), double equidistribution (8.3), and Lemma˜8.10, that
@@
where the second inequality relies on the definition of $J^{\prime}_{2}$. Plugging this bound and Lemma˜8.11 into (8.21), we obtain
@@
We finally deal with $J_{3}$. Applying the Cauchy-Schwarz inequality then Lemma˜8.11, we obtain
@@
We also need the next lemma, which is a variant of [25, Lemma 1.5]. It converts a variance control as in Proposition˜8.12 into an asymptotic estimate.
@@
Let $(Y_{j})_{j\geq 1}$ be a sequence of non-negative real random variables. Let $(y_{j})_{j\geq 1},(y^{\prime}_{j})_{j\geq 1}\in\mathbb{R}_{\geq 0}^{\mathbb{N}}$ be sequences of non-negative real numbers. Set $Z_{j}=Y_{j}-y_{j}$. Assume $y_{j}\leq y_{j}^{\prime}$ for all $j$, as well as $\sum_{j=1}^{\infty}y_{j}=+\infty$, and that for some $C_{1}\geq 1$, for all $n\geq m\geq C_{1}$,
(8.22) $\mathbb{E}\Bigl{[}\bigl{(}\sum_{j=m}^{n}Z_{j}\bigr{)}^{2}\Bigr{]}\leq C_{1}\bigl{(}1+\sum_{j=m}^{n}y^{\prime}_{k}\bigr{)}^{3/2}.$
@@
For an interval $J\subseteq\mathbb{R}_{>0}$, we use the notation $Z_{J}=\sum_{j\in\mathbb{N}_{\geq 1}\cap J}Z_{j}$ and define similarly $y_{J}$, $y^{\prime}_{J}$. We prove the following slightly stronger statement : there is an almost-surely finite random variable $C_{2}$ such that for all $N\geq C_{2}$, we have
(8.23) $\lvert Z_{(0,N]}\rvert\leq(\log(y^{\prime}_{(0,N]}+2))^{2}(y^{\prime}_{(0,N]}+2)^{3/4}+C_{2}.$
For this, up to throwing away a finite number of terms, we may assume (8.22) holds for all $n\geq m\geq 1$.
@@
(8.24) $\forall i\geq 0,\quad y^{\prime}_{(N_{i},N_{i+1}]}\geq 1.$
@@
(8.25) $Z_{(0,N_{i}]}^{2}\leq(\log y^{\prime}_{(0,N_{i}]})^{4}(y^{\prime}_{(0,N_{i}]})^{3/2}.$
Proof of Lemma˜8.14.
@@
The advantage of using $\mathscr{N}$} and not $\mathscr{M}$} to define $N_{i}$ is that we can guarantee (8.24). Thus we can apply Lemma˜8.14 to the sequence $(N_{i})$. We obtain that almost surely, if $N\geq 1$ is sufficiently large, then the unique $i\geq 1$ such that $N\in[N_{i},N_{i+1})$ satisfies (8.25). Recalling that the $Y_{j}$’s are almost-surely non-negative, we obtain that
@@
and the desired lower bound follows, noting that by construction, we have $0\leq y^{\prime}_{(N_{i},N]}\leq 2$.
@@
Applying Lemma˜8.14 with $(N^{\prime}_{i})_{i}$, we have that almost-surely, for large enough $N$, for $i$ such that $(N^{\prime}_{i-1},N^{\prime}_{i}]$,
$Z_{(0,N]}\leq Z_{(0,N^{\prime}_{i}]}+y_{(N,N^{\prime}_{i}]}\leq(\log y^{\prime}_{(0,N^{\prime}_{i}]})^{2}(y^{\prime}_{(0,N^{\prime}_{i}]})^{3/4}+y_{(N,N^{\prime}_{i}]}.$
@@
Combining Proposition˜8.12 and Lemma˜8.13, we obtain the following counting estimate for parameters in $\operatorname{K_{small}}$. Given $n\geq 1$, we recall $\operatorname{K_{small}}(n)=\operatorname{K_{small}}\cap\llbracket 1,n\rrbracket$.
@@
Assume $\sum_{k\in\operatorname{K_{small}}}\psi(\tau^{k})^{d}\tau^{k}=+\infty$ and (8.16), as well as $\gamma_{1}\lll 1$. Then for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}^{d}$, for large enough $n$,
(8.27) $\sum_{k\in\operatorname{K_{small}}(n)}\mathscr{S}_{k}(\boldsymbol{s})\geq(1-\tau^{-1}-C(d)\varepsilon)\sum_{k\in\operatorname{K_{small}}(n)}\psi(\tau^{k})^{d}\tau^{k}$
@@
By the convergent case of the Khintchine dichotomy, replacing $\psi(q)$ by $\max(\psi(q),q^{-1/d}\log^{-2/d}(q))$ may only perturb both sums in (8.27) by a bounded additive constant (which may depend on $\boldsymbol{s}$). As the right hand side of (8.27) is divergent by hypothesis, this perturbation does not affect asymptotics, so we may assume $\psi(q)\geq q^{-1/d}\log^{-2/d}(q)$ for all $q\in\operatorname{K_{small}}$. Recalling (8.8) and $\widetilde{{\mathbbm{1}}}_{R_{k}}\geq\varphi_{k}$, then combining Proposition˜8.12 and Lemma˜8.13, we obtain that $\xi$-almost surely, for large enough $n$,
@@
Proof of Proposition˜8.4.
It follows by combining Lemma˜8.7 and Corollary˜8.15. ∎
@@ 8.2. Case $d\geq 2$, the upper bound
The proof of the asymptotic upper bound in ˜8.3 (case $d\geq 2$) is similar to that of the lower bound. We briefly sketch the proof to highlight the relevant changes. We extend $\psi$ to $\mathbb{R}_{\geq 0}$ by setting $\psi(q)=\psi(\lfloor q\rfloor)$ for non-integer values of $q$. We let $\tau\in(1,2]$ and note that for all integers $N,n\geq 1$ such that $\tau^{n}\leq N<\tau^{n+1}$, for every $\boldsymbol{s}\in\mathbb{R}^{d}$, we have
@@
Then it suffices to show the upper bound analogue of Proposition˜8.4.
@@
Under the assumptions of ˜8.3 and with $d\geq 2$, we have for every $\tau\in(1,2]$, for $\xi$-almost all $\boldsymbol{s}\in\mathbb{R}^{d}$, for every $\eta>0$, for all large enough $n$,
@@
To show Lemma˜8.16, we note that
@@
where $P_{k}:=[0,r_{k})^{d}\times[r_{k},\tau r_{k})$ and $r_{k},t_{k}>0$ have been defined in (8.7). Keep the notations $\gamma_{1}$, $\operatorname{K_{big}}$, $\operatorname{K_{small}}$, from the proof of the lower bound. We establish the upper bound announced in Lemma˜8.16 along the subsums $\sum_{k\in\operatorname{K_{big}}(n)}\mathscr{S}^{+}_{k}(\boldsymbol{s})$ and $\sum_{k\in\operatorname{K_{big}}(n)}\mathscr{S}^{+}_{k}(\boldsymbol{s})$ separately, and under the assumptions that the corresponding $\sum_{k\in\operatorname{K_{big}}}\psi(\tau^{k})^{d}\tau^{k}$, $\sum_{k\in\operatorname{K_{small}}}\psi(\tau^{k})^{d}\tau^{k}$ diverge. This is enough in view of the convergent case of the Khintchine dichotomy, ˜8.2.
The proof of the asymptotic upper bound for $\sum_{k\in\operatorname{K_{big}}(n)}\mathscr{S}^{+}_{k}(\boldsymbol{s})$ is the same as that of Lemma˜8.7, but using this time the upper bound from Lemma˜8.6 instead of the lower bound.
@@
where $\chi_{k}$ is defined as in (8.9). In view of (4.3), $\varphi_{k}^{+}$ coincides with $\theta*\widetilde{{\mathbbm{1}}}_{P_{k}^{+}}$ on the subset $\{\operatorname{dist}(\cdot,x_{0})\leq\frac{\gamma_{2}}{M}\log t_{k}-M\}$ for some $M=M(d)>1$.
By Lemma˜8.5, for $\xi$-almost every $\boldsymbol{s}$, for large enough $k$, we have
$\operatorname{dist}(a(t_{k})u(\boldsymbol{s})x_{0},x_{0})\leq\log\log t_{k},$
and in particular,
$\mathscr{S}^{+}_{k}(\boldsymbol{s})\leq(\theta*\widetilde{{\mathbbm{1}}}_{P_{k}^{+}})(a(t_{k})u(\boldsymbol{s})x_{0})=\varphi^{+}_{k}(a(t_{k})u(\boldsymbol{s})x_{0}).$
Therefore, we only need to show the upper bound analogue of Corollary˜8.15:
(8.28) $\sum_{k\in\operatorname{K_{small}}(n)}\varphi^{+}_{k}(a(t_{k})u(\boldsymbol{s})x_{0})\leq(\tau-1+C(d)\varepsilon)\sum_{k\in\operatorname{K_{small}}(n)}\psi(\tau^{k})^{d}\tau^{k}.$
@@
We have thus established Lemma˜8.16, whence the asymptotic upper bound
@@
It remains to establish ˜8.3 in the case where $d=1$. The proof is similar to the higher dimensional case but a certain number of refinements are required due to poorer moment estimates for Siegel transforms.
Let us start with the lower bound. Keep the notations $\tau$, $R_{k}$, $\gamma_{1}$, $\operatorname{K_{big}}$, $\operatorname{K_{small}}$, $\gamma_{2}$, $\chi_{k}$, $\varepsilon$, $R_{k}^{-}$, $\theta$, from §8.1. The asymptotic lower bound for $\sum_{k\in\operatorname{K_{big}}(n)}\mathscr{S}_{k}(\boldsymbol{s})$ given in Lemma˜8.7 is still valid because the argument works without restriction on $d$.
We thus focus on the lower asymptotic for $\sum_{k\in\operatorname{K_{small}}(n)}\mathscr{S}_{k}(\boldsymbol{s})$, more precisely on extending Corollary˜8.15 to the case $d=1$. The difference with the higher dimensional case is that the Siegel transform of a ball in $\mathbb{R}^{2}$ does not have finite second moment, in particular (8.11) is not valid anymore. To deal with this obstacle, we restrict the Siegel transform by counting only lattice points $(p,q)$ which are bounded multiple of a primitive point. Namely, given $m>0$, we set
@@
where $\operatorname{GCD}(p,q)\in\mathbb{N}_{\geq 1}$ denotes the greatest common divisor of $p$ and $q$. Given a measurable function $f:\mathbb{R}^{2}\rightarrow\mathbb{R}_{\geq 0}$, we define its restricted Siegel transform $\widetilde{f}^{(m)}:X\rightarrow[0,+\infty]$ by: $\forall g\in G$,
@@
In this context, we have the following moment estimates. Their vocation is to replace ˜8.8 from the higher dimensional case.
@@
Let $m\in\mathbb{N}_{\geq 1}$, let $k\in\operatorname{K_{small}}$. Let $c_{m}:=\zeta(2)^{-1}\sum_{t=1}^{m}t^{-2}$ and let $R\subseteq\mathbb{R}^{2}$ be a rectangle such as $R_{k}^{-}$ here and $P_{k}^{+}$ further below. We have
(8.29) $\int_{X}\widetilde{{\mathbbm{1}}}^{(m)}_{R}\,\mathrm{d}m_{X}=c_{m}\operatorname{Leb}(R)$
(8.30) $\int_{X}(\widetilde{{\mathbbm{1}}}^{(m)}_{R})^{2}\,\mathrm{d}m_{X}=c_{m}^{2}\operatorname{Leb}(R)^{2}+O\left(\operatorname{Leb}(R)\log(1+m)\right).$
@@
Those estimates appear in the litterature for primitive Siegel transforms, i.e. for $m=1$. We explain how to deduce the case $m\geq 1$. Note that
$\widetilde{{\mathbbm{1}}}^{(m)}_{R}=\sum_{q\in\llbracket 1,m\rrbracket}\widetilde{{\mathbbm{1}}}^{(1)}_{R/q}$.
The Siegel summation formula for primitive Siegel transform [50, Equation ($8$)] guarantees $\int_{X}\widetilde{{\mathbbm{1}}}^{(1)}_{R/q}\,\mathrm{d}m_{X}=\zeta(2)^{-1}\operatorname{Leb}(R/q)$, whence (8.29).
To justify (8.30), we note that
$\int_{X}(\widetilde{{\mathbbm{1}}}^{(m)}_{R})^{2}\,\mathrm{d}m_{X}=\sum_{q,q^{\prime}\in\llbracket 1,m\rrbracket}\int_{X}\widetilde{{\mathbbm{1}}}^{(1)}_{R/q}\widetilde{{\mathbbm{1}}}^{(1)}_{R/q^{\prime}}\,\mathrm{d}m_{X}.$
By Rogers [48, Theorem 5], we have
$\int_{X}\widetilde{{\mathbbm{1}}}^{(1)}_{R/q}\widetilde{{\mathbbm{1}}}^{(1)}_{R/q^{\prime}}\,\mathrm{d}m_{X}$
$=\zeta(2)^{-2}q^{-2}q^{\prime-2}\operatorname{Leb}(R)^{2}+\zeta(2)^{-1}\operatorname{Leb}(R/q\cap R/q^{\prime})+\zeta(2)^{-1}\operatorname{Leb}(R/q\cap(-R/q^{\prime})).$
Summing over $q,q^{\prime}\in\llbracket 1,m\rrbracket$ the first term of the right hand side gives
$\sum_{q,q^{\prime}\in\llbracket 1,m\rrbracket}\zeta(2)^{-2}q^{-2}q^{\prime-2}\operatorname{Leb}(R)^{2}=c_{m}^{2}\operatorname{Leb}(R)^{2}.$
On the other hand, looking at the second coordinate of points in $R$, we find for any $q\geq 1$,
$|\{q^{\prime}\geq 1\,:\,R/q\cap R/q^{\prime}\neq\varnothing\}|\ll q$
Using the trivial bound $\operatorname{Leb}(R/q\cap R/q^{\prime})\leq\operatorname{Leb}(R)/q^{2}$, we deduce
$\sum_{q,q^{\prime}\in\llbracket 1,m\rrbracket}\operatorname{Leb}(R/q\cap R/q^{\prime})\ll\sum_{q\in\llbracket 1,m\rrbracket}\operatorname{Leb}(R)/q\simeq\operatorname{Leb}(R)\log(1+m).$
As $R/q\cap(-R/q^{\prime})=\varnothing$, the last term does not contribute, and this concludes the proof of (8.30). ∎
@@
(8.31) $\forall k\in\operatorname{K_{small}},\quad\log(1+m_{k})=\Bigl{(}\sum_{j\in\operatorname{K_{small}}(k)}\psi(\tau^{j})\tau^{j}\Bigr{)}^{1/8}.$

### v2

@@ 9.1. Case $d\geq 2$, the lower bound
$\tau^{-cj/2}\ll(j\log\tau)^{-2}\leq r_{j}^{d+1}\ll y_{j},$
where the last inequality relies on the estimate (9.14), and the bound $t_{j}^{-\gamma_{2}/M}\leq(j\log\tau)^{-2}\leq r_{j}^{d+1}$ (which holds under the assumption $\inf J\ggg_{\gamma_{2}}1$). Hence $\tau^{-ck/2}\ll y_{j}\tau^{-c(k-j)/2}$, so $\sum\nolimits_{j,k\in J_{1},j\leq k}\tau^{-ck/2}\ll y_{J_{1}}$ as for the first sum. The fourth sum can be handled similarly to the third one. In the end, we have bounded $\mathbb{E}[Z_{J_{1}}^{2}]$ by $O(\sum_{J_{1}}y_{j})\ll(1+y_{J_{1}})^{3/2}$ as desired.
@@
Set $m:=|J_{2}|$. We start with the case where $m$ is very small compared to $n=|J_{2}\sqcup J_{3}|$, more precisely we assume $m^{2}\leq n$. In this scenario, we have by the Cauchy-Schwarz inequality and Lemma˜9.11,
@@
$\mathbb{E}\bigl[Z_{J^{\prime}_{2}}^{2}\bigr]$ $=\sum_{|j-k|<\sqrt{m}}\mathbb{E}[Z_{j}Z_{k}]+\sum_{|j-k|\geq\sqrt{m}}\mathbb{E}[Z_{j}Z_{k}]$ (9.22)
$\ll\sqrt{m}\sum_{j\in J^{\prime}_{2}}\mathbb{E}[Z_{j}^{2}]+\Bigl\lvert\sum_{|j-k|\geq\sqrt{m}}\mathbb{E}[Z_{j}Z_{k}]\Bigr\rvert$ (9.23)
@@
For subscripts $j\leq k\in J_{2}^{\prime}$ such that $|j-k|\geq\sqrt{m}$, we have by (9.19), double equidistribution (9.3), and Lemma˜9.10, that
@@
where the second inequality relies on the definition of $J^{\prime}_{2}$. Plugging this bound and Lemma˜9.11 into (9.23), we obtain
@@
We finally deal with $J_{3}$. Applying the Cauchy-Schwarz inequality then Lemma˜9.11, we obtain
@@
We also need the next lemma, which is a variant of [28, Lemma 1.5]. It converts a variance control as in Proposition˜9.12 into an asymptotic estimate.
@@
Let $(Y_{j})_{j\geq 1}$ be a sequence of non-negative real random variables. Let $(y_{j})_{j\geq 1},(y^{\prime}_{j})_{j\geq 1}\in\mathbb{R}_{\geq 0}^{\mathbb{N}}$ be sequences of non-negative real numbers. Set $Z_{j}=Y_{j}-y_{j}$. Assume $\sum_{j=1}^{\infty}y_{j}=+\infty$, and that for some $C_{1}\geq 1$, for all $n\geq m\geq C_{1}$, we have $y_{m}\leq y_{m}^{\prime}$ and
$\mathbb{E}\Bigl[\bigl(\sum_{j=m}^{n}Z_{j}\bigr)^{2}\Bigr]\leq C_{1}\bigl(1+\sum_{j=m}^{n}y^{\prime}_{k}\bigr)^{3/2}.$ (9.24)
@@
For an interval $J\subseteq\mathbb{R}_{>0}$, we use the notation $Z_{J}=\sum_{j\in\mathbb{N}\cap J}Z_{j}$ and define similarly $y_{J}$, $y^{\prime}_{J}$. We prove the following slightly stronger statement : there is an almost-surely finite random variable $C_{2}$ such that for all $N\geq C_{2}$, we have
$\lvert Z_{(0,N]}\rvert\leq(\log(y^{\prime}_{(0,N]}+2))^{2}(y^{\prime}_{(0,N]}+2)^{3/4}+C_{2}.$ (9.25)
For this, up to throwing away a finite number of terms, we may assume that the relations $y_{m}\leq y^{\prime}_{m}$ and (9.24) hold for all $n\geq m\geq 1$.
@@
$\forall i\geq 0,\qquad y^{\prime}_{(N_{i},N_{i+1}]}\geq 1.$ (9.26)
@@
$Z_{(0,N_{i}]}^{2}\leq(\log y^{\prime}_{(0,N_{i}]})^{4}(y^{\prime}_{(0,N_{i}]})^{3/2}.$ (9.27)
Proof of Lemma˜9.14.
@@
The advantage of using $\mathscr{N}$} and not $\mathscr{M}$} to define $N_{i}$ is that we can guarantee (9.26). Thus we can apply Lemma˜9.14 to the sequence $(N_{i})$. We obtain that almost surely, if $N\geq 1$ is sufficiently large, then the unique $i\geq 1$ such that $N\in[N_{i},N_{i+1})$ satisfies (9.27). Recalling that the $Y_{j}$’s are almost-surely non-negative, we obtain that
@@
and the desired lower bound follows as noting that by construction, we have $0\leq y^{\prime}_{(N_{i},N]}\leq 2$.
@@
Applying Lemma˜9.14 with $(N^{\prime}_{i})_{i}$, we have that almost-surely, for large enough $N$, for $i$ such that $N\in(N^{\prime}_{i-1},N^{\prime}_{i}]$,
$Z_{(0,N]}\leq Z_{(0,N^{\prime}_{i-1}]}+y_{(N,N^{\prime}_{i}]}\leq(\log y^{\prime}_{(0,N^{\prime}_{i}]})^{2}(y^{\prime}_{(0,N^{\prime}_{i}]})^{3/4}+y_{(N,N^{\prime}_{i}]}.$
@@
Combining Proposition˜9.12 and Lemma˜9.13, we obtain the following counting estimate for parameters in $\operatorname{K_{small}}$. Given $n\geq 1$, we recall $\operatorname{K_{small}}(n)=\operatorname{K_{small}}\cap\llbracket 1,n\rrbracket$.
@@
Assume $\sum_{k\in\operatorname{K_{small}}}\psi(\tau^{k})^{d}\tau^{k}=+\infty$, as well as $\gamma_{1}\lll 1$. Then for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}^{d}$, for large enough $n$,
$\sum_{k\in\operatorname{K_{small}}(n)}\mathscr{S}_{k}(\boldsymbol{s})\geq(1-\tau^{-1}-C(d)\varepsilon)\sum_{k\in\operatorname{K_{small}}(n)}\psi(\tau^{k})^{d}\tau^{k}$ (9.29)
@@
By the convergent case of the Khintchine dichotomy, replacing $\psi(q)$ by $\max(\psi(q),q^{-1/d}\log^{-2/d}(q))$ may only perturb both sums in (9.29) by a bounded additive constant (which may depend on $\boldsymbol{s}$). As the right hand side of (9.29) is divergent by hypothesis, this perturbation does not affect asymptotics, so we may assume $\psi(q)\geq q^{-1/d}\log^{-2/d}(q)$ for all $q\in\operatorname{K_{small}}$. Recalling (9.10) and $\widetilde{{\mathbbm{1}}}_{R_{k}}\geq\varphi_{k}$, then combining Proposition˜9.12 and Lemma˜9.13, we obtain that $\xi$-almost surely, for large enough $n$,
@@
Proof of Proposition˜9.4.
It follows by combining Lemma˜9.7 and Corollary˜9.15. ∎
@@ 9.2. Case $d\geq 2$, the upper bound
The proof of the asymptotic upper bound in ˜9.3 (case $d\geq 2$) is similar to that of the lower bound. We briefly sketch the proof to highlight the relevant changes. We extend $\psi$ to $\mathbb{R}_{\geq 0}$ by setting $\psi(q)=\psi(\lfloor q\rfloor)$ for non-integer values of $q$. We let $\tau\in(1,2]$ and note that for all integers $N,n\geq 1$ such that $\tau^{n}\leq N<\tau^{n+1}$, for every $\boldsymbol{s}\in\mathbb{R}^{d}$, we have
@@
Then it suffices to show the upper bound analogue of Proposition˜9.4.
@@
Under the assumptions of ˜9.3 and with $d\geq 2$, we have for every $\tau\in(1,2]$, for $\xi$-almost all $\boldsymbol{s}\in\mathbb{R}^{d}$, for every $\eta>0$, for all large enough $n$,
@@
To show Lemma˜9.16, we note that
@@
where $P_{k}:=[0,r_{k})^{d}\times[r_{k},\tau r_{k})$ and $r_{k},t_{k}>0$ have been defined in (9.9). Keep the notations $\gamma_{1}$, $\operatorname{K_{big}}$, $\operatorname{K_{small}}$, from the proof of the lower bound. We establish the upper bound announced in Lemma˜9.16 along the subsums $\sum_{k\in\operatorname{K_{big}}(n)}\mathscr{S}^{+}_{k}(\boldsymbol{s})$ and $\sum_{k\in\operatorname{K_{small}}(n)}\mathscr{S}^{+}_{k}(\boldsymbol{s})$ separately, and under the assumptions that the corresponding $\sum_{k\in\operatorname{K_{big}}}\psi(\tau^{k})^{d}\tau^{k}$, $\sum_{k\in\operatorname{K_{small}}}\psi(\tau^{k})^{d}\tau^{k}$ diverge. This is enough in view of the convergent case of the Khintchine dichotomy, ˜9.2.
The proof of the asymptotic upper bound for $\sum_{k\in\operatorname{K_{big}}(n)}\mathscr{S}^{+}_{k}(\boldsymbol{s})$ is the same as that of Lemma˜9.7, but using this time the upper bound from Lemma˜9.6 instead of the lower bound.
@@
where $\chi_{k}$ is defined as in (9.11). Note that contrary to the lower bound case, the truncation plays against comparing $\widetilde{{\mathbbm{1}}}_{P_{k}}$ with $\varphi_{k}^{+}$. However, in view of (5.3), $\varphi_{k}^{+}$ coincides with $\theta*\widetilde{{\mathbbm{1}}}_{P_{k}^{+}}$ on the subset $\{\operatorname{dist}(\cdot,x_{0})\leq\frac{\gamma_{2}}{M}\log t_{k}-M\}$ for some $M=M(d)>1$, in particular we do have $\widetilde{{\mathbbm{1}}}_{P_{k}}\leq\varphi_{k}^{+}$ on that domain. This is enough for us because, for $\xi$-typical $\boldsymbol{s}$, and large enough $k$, the points $a(t_{k})u(\boldsymbol{s})x_{0}$ are in this domain, indeed Lemma˜9.5 gives
We get for $\xi$-almost every $\boldsymbol{s}$, for large enough $k$,
$\mathscr{S}^{+}_{k}(\boldsymbol{s})\leq\varphi^{+}_{k}(a(t_{k})u(\boldsymbol{s})x_{0}).$
Therefore, we only need to show the upper bound analogue of Corollary˜9.15:
$\sum_{k\in\operatorname{K_{small}}(n)}\varphi^{+}_{k}(a(t_{k})u(\boldsymbol{s})x_{0})\leq(\tau-1+C(d)\varepsilon)\sum_{k\in\operatorname{K_{small}}(n)}\psi(\tau^{k})^{d}\tau^{k}$ (9.30)
@@
We have thus established Lemma˜9.16, whence the asymptotic upper bound
@@
It remains to establish ˜9.3 in the case where $d=1$. The proof is similar to the higher dimensional case but a certain number of refinements are required due to poorer moment estimates for Siegel transforms.
Let us start with the lower bound. Keep the notations $\tau$, $R_{k}$, $\gamma_{1}$, $\operatorname{K_{big}}$, $\operatorname{K_{small}}$, $\gamma_{2}$, $\chi_{k}$, $\varepsilon$, $R_{k}^{-}$, $\theta$, from §9.1. The asymptotic lower bound for $\sum_{k\in\operatorname{K_{big}}(n)}\mathscr{S}_{k}(\boldsymbol{s})$ given in Lemma˜9.7 is still valid because the argument works without restriction on $d$.
We thus focus on the lower asymptotic for $\sum_{k\in\operatorname{K_{small}}(n)}\mathscr{S}_{k}(\boldsymbol{s})$, more precisely on extending Corollary˜9.15 to the case $d=1$. The difference with the higher dimensional case is that the Siegel transform of a ball in $\mathbb{R}^{2}$ does not have finite second moment, in particular (9.13) is not valid anymore. To deal with this obstacle, we restrict the Siegel transform by counting only lattice points $(p,q)$ which are bounded multiple of a primitive point. Namely, given $m>0$, we set
@@
where $\operatorname{GCD}(p,q)\in\mathbb{N}^{*}$ denotes the greatest common divisor of $p$ and $q$. Given a measurable function $f:\mathbb{R}^{2}\rightarrow\mathbb{R}_{\geq 0}$, we define its restricted Siegel transform $\widetilde{f}^{(m)}:X\rightarrow[0,+\infty]$ by: $\forall g\in G$,
@@
In this context, we have the following moment estimates. Their vocation is to replace ˜9.8 from the higher dimensional case.
@@
Let $m\in\mathbb{N}^{*}$, let $k\in\operatorname{K_{small}}$. Let $c_{m}:=\zeta(2)^{-1}\sum_{t=1}^{m}t^{-2}$ and let $R\subseteq\mathbb{R}^{2}$ be a bounded measurable subset. We have
$\int_{X}\widetilde{{\mathbbm{1}}}^{(m)}_{R}\,\mathrm{d}m_{X}=c_{m}\operatorname{Leb}(R)$ (9.31)
$\int_{X}(\widetilde{{\mathbbm{1}}}^{(m)}_{R})^{2}\,\mathrm{d}m_{X}=c_{m}^{2}\operatorname{Leb}(R)^{2}+O\left(\operatorname{Leb}(R)\log^{2}(1+m)\right).$ (9.32)
@@
Those estimates appear in the literature for primitive Siegel transforms, i.e., for $m=1$. We explain how to deduce the case $m\geq 1$. Note that
$\widetilde{{\mathbbm{1}}}^{(m)}_{R}=\sum_{q\in\llbracket 1,m\rrbracket}\widetilde{{\mathbbm{1}}}^{(1)}_{R/q}=\widetilde{h}^{(1)}$
where $h:=\sum_{q\in\llbracket 1,m\rrbracket}{{\mathbbm{1}}}_{R/q}$. The Siegel summation formula for primitive Siegel transform [58, Equation (88)] guarantees $\int_{X}\widetilde{h}^{(1)}\,\mathrm{d}m_{X}=\zeta(2)^{-1}\int_{\mathbb{R}^{2}}h\,\,\mathrm{d}\text{Leb}=c_{m}\operatorname{Leb}(R)$, whence (9.31).
To justify (9.32), we apply [34, Theorem 1 & Remark 0.9], and the Cauchy-Schwarz inequality to bound the integral involving the isometry in that theorem, to get
$\int_{X}(\widetilde{h}^{(1)})^{2}\,\mathrm{d}m_{X}$ $=\left|\zeta(2)^{-1}\int_{\mathbb{R}^{2}}h\,\,\mathrm{d}\text{Leb}\right|^{2}+O\left(\int_{\mathbb{R}^{2}}h^{2}\,\,\mathrm{d}\text{Leb}\right).$
The first term on the right-hand side is equal to $c_{m}^{2}\operatorname{Leb}(R)^{2}$. The second term can be bounded via
$\int_{\mathbb{R}^{2}}h^{2}\,\,\mathrm{d}\text{Leb}$ $=\sum_{q,q^{\prime}\in\llbracket 1,m\rrbracket}\int_{\mathbb{R}^{2}}{\mathbbm{1}}_{R/q}{\mathbbm{1}}_{R/q^{\prime}}\,\,\mathrm{d}\text{Leb}\leq\sum_{q,q^{\prime}\in\llbracket 1,m\rrbracket}\|{\mathbbm{1}}_{R/q}\|_{L^{2}}\|{\mathbbm{1}}_{R/q^{\prime}}\|_{L^{2}}$
$\leq\operatorname{Leb}(R)\sum_{q,q^{\prime}\in\llbracket 1,m\rrbracket}q^{-1}q^{\prime-1}\ll\operatorname{Leb}(R)\log^{2}(1+m).$
This concludes the proof of (9.32). ∎
@@
$\forall k\in\operatorname{K_{small}},\qquad\log^{2}(1+m_{k})=\Bigl(\sum_{j\in\operatorname{K_{small}}(k)}\psi(\tau^{j})\tau^{j}\Bigr)^{1/8}.$ (9.33)